# Trilinos: Enabling Scientific Computing Across Diverse Hardware Architectures at Scale

### v1

@@ 2.6. Combinatorial algorithms with hybrid parallelism: Zoltan2
Zoltan2 is a package for partitioning, load balancing, and other combinatorial scientific computing. It can be viewed as a successor to Zoltan (Boman et al., 2012) that provides both tighter integration with other Trilinos packages and support for hybrid parallelism. It is used primarily to partition and distribute data objects (such as sparse matrices) across processors.
@@
Zoltan2 supports both distributed-memory parallelism and on-node parallelism for both CPU and GPU (via Kokkos). Algorithms that support such hybrid parallelism include Multi-jagged (Deveci et al., 2016) (geometric partitioning), Sphynx (Acer et al., 2021) (spectral graph partitioning), graph coloring (Bogle et al., 2020), and MPI task placement (Deveci et al., 2019).
@@ 3.1. Iterative Linear Solvers and Krylov Methods: Belos
Belos (Bavier et al., 2012) provides next-generation iterative linear solvers and a powerful developer framework for solving linear systems and least-squares problems. This framework provides several abstractions that facilitate code reuse and extensibility.
First, Belos decouples the algorithms from the implementation of the underlying linear algebra object using traits mechanisms. While concrete linear algebra adapters are provided for Tpetra and Thyra, users can also implement their own interfaces to leverage any existing matrix and vector representations. Second, there are abstractions to orthogonalization to ease the integration of application or architecture-specific orthogonalization methods. Implementations of iterated classical Gram-Schmidt, DGKS-corrected (Daniel, Gragg, Kaufman, and Stewart-corrected) classical Gram-Schmidt, and iterated modified Gram-Schmidt algorithms are provided. Finally, powerful solver managers encapsulate the strategy for solving a linear system or least-squares problem using abstract interfaces to iteration kernels. As a result, the algorithms developed using Belos abstractions can be relatively agnostic to data layout in memory or distributed over processors and parallel matrix/vector operations.
Belos supersedes the AztecOO package (Heroux, 2004) providing solver managers for single-vector iterative solvers, but also extensions of these methods to block iterative methods for solving linear systems with multiple right-hand sides. Single-vector solver managers include conjugate gradient (CG), minimum residual (MINRES), generalized minimal residual (GMRES), stabilized biconjugate gradient (BiCGStab), and transpose-free QMR (TFQMR) as well as “seed” solvers (hybrid GMRES, PCPG), subspace recycling solvers (GCRO-DR and RCG), and least-squares solvers (LSQR). There are versions of CG, GMRES, and GCRO-DR that construct a block Krylov subspace to solve for one or more right-hand sides. There are also “pseudo-block” versions of CG, GMRES, and TFQMR that apply the single-vector iteration simultaneously on multiple right-hand sides, where matrix and preconditioner applications are aggregated to achieve better performance. Any single-vector iteration can be used to solve multiple right-hand sides as well but, if a block or pseudo-block version is not used, the solver will sequentially solve the right-hand sides.
@@ 3.2. One-Level Domain Decomposition and Basic Iterative Methods: Ifpack2
[…]
@@ 3.3. Multilevel Domain Decomposition Methods: FROSch
FROSch (Fast and Robust Overlapping Schwarz) is a framework for the construction of multilevel Schwarz domain decomposition preconditioners. Besides parallel scalability, FROSch emphasizes applicability and robustness across a wide range of challenging problems while supporting an algebraic construction. Specifically, most preconditioners can be built using only the fully assembled system matrix, though some variants can take advantage of additional geometric inputs. The algebraic construction relies on two key algorithmic components: first, the creation of an overlapping domain decomposition at the initial level based on the sparsity pattern of the system matrix, similar to Ifpack2; and second, the integration of extension-based coarse spaces, as in the classical two-level Generalized Dryja–Smith–Widlund (GDSW) preconditioner (Dohrmann et al., 2008) and its variants. While the initial version of FROSch (Heinlein et al., 2016) was based on the outdated Epetra linear algebra framework, the current implementation (Heinlein et al., 2020) leverages Xpetra. Originally designed as a lightweight wrapper around Epetra and Tpetra, the Xpetra package facilitated over the years compatibility with both the Epetra and Tpetra stacks. With the deprecation of Epetra, it now exclusively supports the Tpetra stack. Algorithmic variants of Schwarz methods implemented in FROSch include:
@@
FROSch has been applied to various challenging application problems, including: scalar elliptic and elasticity problems (Heinlein et al., 2016), possibly with heterogeneities (Alves et al., 2024); computational fluid dynamics problems (Heinlein et al., 2019a, [n. d.]); time-harmonic Maxwell’s and fluid-structure interaction problems (Heinlein et al., 2024); pharmaco-mechanical interactions in arterial walls (Balzani et al., [n. d.]); and coupled multiphysics problems for land ice simulations (Heinlein et al., 2022a). The latter three have been solved using monolithic preconditioning techniques. To extend robustness for heterogeneous model problems, an implementation of spectral coarse spaces (Heinlein et al., 2019b) is currently under development. FROSch preconditioners have scaled to more than $200,000$ cores on the Theta Cray XC40 supercomputer at the Argonne Leadership Computing Facility (ALCF); cf. (Heinlein et al., 2022b).
In its current implementation, FROSch assumes a one-to-one correspondence of subdomains and MPI ranks. Using an interface to the other solver packages in Trilinos, a variety of inexact solvers can be employed for the solution of the local subdomain problems. An extension to multiple subdomains per MPI rank is currently being implemented. Using Kokkos and Kokkos Kernels, FROSch has recently also been ported to GPUs (Yamazaki et al., 2023) with performance gains for the triangular solve or inexact solves with ILU on GPUs.
A demo/tutorial for FROSch can be found at the GitHub repository (fro, [n. d.]).
@@ 3.4. Multigrid Methods: MueLu
MueLu is a flexible and scalable high-performance multigrid solver library. It provides a variety of multigrid algorithms for problems ranging from Poisson-like operators, over elasticity, convection-diffusion, Navier-Stokes, and Maxwell’s equations, all the way to multigrid methods for coupled multiphysics systems. Besides its strong focus on aggregation-based algebraic multigrid (AMG) methods, MueLu comes with specialized capabilities for (semi-)structured grids to perform semi-coarsening along grid lines while forming the coarse operator via a Galerkin product (in contrast to classical geometric multigrid methods). MueLu is extensible and allows for the research and development of new multigrid preconditioning methods. Its weak and strong scalability up to $131,000$ cores of a Cray XC40 and one million cores of a Blue Gene/Q system, even for vector-valued partial differential equations (PDEs) on unstructured meshes, have been shown in (Lin et al., 2017; Thomas et al., 2019).
@@
[…]
@@
There are a variety of resources to learn more about MueLu. An overview is given on the MueLu website [^10] https://trilinos.github.io/muelu.html. The MueLu User’s Guide (Berger-Vergiat et al., 2023) summarizes installation instructions and a reference to most of MueLu’s configuration parameters. The MueLu Tutorial (Mayr et al., 2023) introduces beginners and experts to various topics in MueLu and shows how to solve or precondition different linear systems using MueLu. Details on the compatibility of MueLu and its predecessor ML (Heroux et al., 2005; Gee et al., 2006) can be found in the MueLu User’s Guide (Berger-Vergiat et al., 2023).
Besides its C++ API, MueLu offers a MATLAB interface, MueMex, to provide access to MueLu’s aggregation and solver routines from MATLAB. MueMex allows users to setup and solve arbitrarily many problems with either MueLu as a preconditioner, Belos as a solver and Epetra or Tpetra for data structures.
@@
3.5. Direct Linear Solvers: ShyLU and Amesos2
ShyLU (Thornquist and Rajamanickam, 2014) provides an implementation of three solvers: Basker, Tacho, and a distributed-memory solver based on the Schur complement method.
Basker (Booth et al., 2017) is a sparse direct solver based on LU factorization for the problems that have the block triangular form (BTF) typically seen in circuit simulation applications. Basker uses this structure to factor and solve the diagonal blocks in parallel. The larger diagonal blocks can themselves be factored in parallel by discovering the parallelism available using a nested-dissection reordering. Basker focuses on exploiting thread-parallelism on the multi-core CPU architectures. Amesos2 also has a templated implementation of a sequential KLU solver called KLU2, which also exploits the BTF structure.
Tacho is a sparse direct solver that exploits supernodal block structure commonly found in sparse direct factorizations of matrices from mechanics applications. Tacho exploits this supernodal structure for both factorization and triangular-solve phases. It is based on Kokkos. Originally, Tacho implemented a task-parallel Cholesky factorization of sparse symmetric positive definite (SPD) matrices (Kim et al., 2018). However, to improve its portability, it has been extended to compute the sparse factorization based on level-set scheduling. Moreover, its functionality has been extended to the computation of an LDLT factorization of symmetric indefinite matrices, as well as the computation of an LU factorization of general matrices with a symmetric sparsity pattern.
@@ 3.5. Direct Linear Solvers: ShyLU and Amesos2
Amesos2 (Bavier et al., 2012) provides a uniform interface to direct linear solvers. These include third-party direct solvers such as CHOLMOD, MUMPS, Pardiso_MKL, SuperLU, SuperLU_MT, SuperLU_Dist, and STRUMPACK. Amesos2 also provides an interface to the on-node sparse direct solvers implemented in ShyLU.
@@ 3.6. Physics Block Operators and Preconditioners: Teko
The Teko library (Cyr et al., 2016) provides interfaces for operators and preconditioners that are constructed from large physics-based sub-blocks. The sub-blocks are Thyra operators which are themselves often implemented using Tpetra matrices. Teko provides generic block preconditioning strategies such as $2\times 2$ block-LU factorizations as well as block versions of Jacobi- and Gauss–Seidel-type methods for any number of blocks, as well as commonly used approximate inverse strategies for the Navier-Stokes equation such as SIMPLEC, LSC, and PCD preconditioners (Cyr et al., 2012). The Block-LU class offers the possibility to implement user-specific inverse approximation strategies, for example using a sparse approximate inverse to construct the Schur complement (Firmbach et al., 2024). More complicated multilevel hierarchies of block solvers can be generated via Teuchos::ParameterList objects. Block preconditioners for first-order formulations of Maxwell’s equations and Darcy flow are implemented in the Panzer package. Teko’s solvers can be registered in the Stratimikos interface for usage entirely driven by Teuchos::ParameterList objects.
@@ 3.8. Unified Solver Interface: Stratimikos
The Stratimikos package offers a cohesive interface to various linear solvers and preconditioners within Trilinos, including Amesos2, Belos, FROSch, Ifpack2, MueLu, and Teko. It requires that the matrix, right-hand side, and the solution vectors all support the Thyra interface. Thyra also provides wrappers for Tpetra linear algebra. Users can specify solver and preconditioner parameters via a Teuchos::ParameterList, which can be easily populated from an XML file. Moreover, the package allows for the easy addition of new solver and preconditioner factories via adapters, with examples available in the packages/stratimikos/adapters directory.
@@ 4. Nonlinear Solvers and Analysis Tools
The packages included in this section provide the top-level algorithms for computational simulations and design studies. These include nonlinear solvers, bifurcation tracking, stability analysis, parameter continuation, optimization, and uncertainty quantification. A common theme of this collection is the philosophy of “analysis beyond simulation”, which aims to automate many computational tasks that are often performed by application code users by trial-and-error or repeated simulation. Tasks that can be automated include performing parameter studies, sensitivity analysis, calibration, optimization, and the task of locating instabilities. Additional utilities for nonlinear analysis include abstraction layers and interfaces for application callbacks, as well as automatic differentiation tools that can provide the derivatives critical to analysis algorithms.

### v2

@@ 2.6. Combinatorial algorithms with hybrid parallelism: Zoltan2
Zoltan2 is a package for partitioning, load balancing, and other combinatorial scientific computing algorithms. It can be viewed as a successor to Zoltan (Boman et al., 2012) that provides both tighter integration with other Trilinos packages and support for hybrid parallelism. It is used primarily to partition and distribute data objects (such as sparse matrices) across processors.
@@
Zoltan2 supports the use of both distributed-memory parallelism and on-node parallelism on CPUs and GPUs by building on Kokkos. Algorithms that support such hybrid parallelism include Multi-jagged (Deveci et al., 2016) (geometric partitioning), Sphynx (Acer et al., 2021) (spectral graph partitioning), graph coloring (Bogle et al., 2020), and MPI task placement (Deveci et al., 2019).
@@ 3.1. Iterative Linear Solvers and Krylov Methods: Belos
Belos (Bavier et al., 2012) provides next-generation iterative linear solvers and a powerful framework to develop solvers for linear systems and least-squares problems. This framework is powerful because it provides several algorithmic abstractions that facilitate code reuse and extensibility. These abstractions are at every level of the iterative solver hierarchy, from the foundational linear algebra to the solver managers that embody the overall solution strategy, facilitating flexible solver development.
Belos fundamentally decouples the algorithms from the implementation of the underlying linear algebra object using abstract traits classes that prescribe the required matrix and vector operations. Adapters are necessary to concrete linear algebra implementations, which are provided for Tpetra and Thyra. Users can also implement their own interfaces to leverage existing matrix and vector representations. Next, there are abstractions to orthogonalization to ease the integration of application or architecture-specific orthogonalization methods. Implementations of iterated classical Gram-Schmidt, DGKS-corrected (Daniel, Gragg, Kaufman, and Stewart-corrected) classical Gram-Schmidt, and iterated modified Gram-Schmidt algorithms are provided. Finally, solver managers encapsulate the strategy for solving a linear system or least-squares problem using abstract interfaces to iteration kernels. As a result, the algorithms developed using Belos abstractions can be relatively agnostic to data layout in memory or distributed over processors and parallel matrix/vector operations.
Belos supersedes the AztecOO package (Heroux, 2004) by providing solver managers for individual iterative methods. A solver manager wraps a specific iteration and provides the high-level control logic and supporting functionality for one method, such as CG or GMRES, rather than for a collection of methods. Belos includes solver managers for single-vector iterative solvers, as well as extensions of several of these methods to block iterative methods for solving linear systems with multiple right-hand sides. Single-vector solver managers include conjugate gradient (CG), minimum residual (MINRES), generalized minimal residual (GMRES), stabilized biconjugate gradient (BiCGStab), and transpose-free QMR (TFQMR) as well as “seed” solvers (hybrid GMRES, PCPG), subspace recycling solvers (GCRO-DR and RCG), and least-squares solvers (LSQR). There are versions of CG, GMRES, and GCRO-DR that construct a block Krylov subspace to solve for one or more right-hand sides. There are also “pseudo-block” versions of CG, GMRES, and TFQMR that apply the single-vector iteration simultaneously on multiple right-hand sides, where matrix and preconditioner applications are aggregated to achieve better performance. Any single-vector iteration can be used to solve multiple right-hand sides as well but, if a block or pseudo-block version is not used, the solver will sequentially solve the right-hand sides.
@@ 3.2. One-Level Domain Decomposition and Basic Iterative Methods: Ifpack2
Ifpack2 also supplies classic iterative methods based on matrix-splitting techniques, such as Jacobi iteration, Gauss-Seidel, and an MPI-oriented hybrid of Jacobi and Gauss-Seidel (e.g., Jacobi between ranks and Gauss-Seidel on them). Ifpack2 also provides preconditioners based on Chebyshev iterations. The aforementioned preconditioners are available both in point and block forms and can operate on CRS and BSR matrices. In the block case, line relaxation is also supported, while in the point case, techniques like Vanka relaxation (Vanka, 1986) are possible. Auxiliary-space smoothing for Maxwell and Darcy type problems in the style of Hiptmair (Hiptmair, 1997) are also supported. Local kernels are either implemented in Ifpack2 itself but can also be called from Kokkos Kernels for performance-portable shared-memory algorithmic variants of Gauss-Seidel and Jacobi iterations.
@@ 3.3. Multilevel Domain Decomposition Methods: FROSch
FROSch (Fast and Robust Overlapping Schwarz) is a framework for the construction of multilevel Schwarz domain decomposition preconditioners. Besides parallel scalability, FROSch emphasizes applicability and robustness across a wide range of challenging problems while supporting algebraic construction of the preconditioners, that is, solely from the fully assembled system matrix.
Most FROSch preconditioners can be constructed algebraically, though some variants can take advantage of additional geometric input. The algebraic construction relies on two key algorithmic components: first, the creation of an overlapping domain decomposition at the initial level based on the sparsity pattern of the system matrix, similar to Ifpack2; and second, the integration of extension-based coarse spaces, as in the classical two-level Generalized Dryja–Smith–Widlund (GDSW) preconditioner (Dohrmann et al., 2008) and its variants. While the initial version of FROSch (Heinlein et al., 2016) was based on the outdated Epetra linear algebra framework, the current implementation (Heinlein et al., 2020) leverages Xpetra. Originally designed as a lightweight wrapper around Epetra and Tpetra, the Xpetra package facilitated over the years compatibility with both the Epetra and Tpetra stacks. With the deprecation of Epetra, it now exclusively supports the Tpetra stack. Algorithmic variants of Schwarz methods implemented in FROSch include:
@@
The wide applicability of FROSch has been demonstrated for various challenging application problems, including: scalar elliptic and elasticity problems (Heinlein et al., 2016), possibly with heterogeneities (Alves et al., 2024); computational fluid dynamics problems (Heinlein et al., 2019a, ); time-harmonic Maxwell’s and fluid-structure interaction problems (Heinlein et al., 2024); pharmaco-mechanical interactions in arterial walls (Balzani et al., ); and coupled multiphysics problems for land ice simulations (Heinlein et al., 2022a). The latter three have been solved using monolithic preconditioning techniques. To extend robustness for heterogeneous model problems, an implementation of spectral coarse spaces (Heinlein et al., 2019b) is currently under development. FROSch preconditioners have scaled to more than $200,000$ cores on the Theta Cray XC40 supercomputer at the Argonne Leadership Computing Facility (ALCF); cf. (Heinlein et al., 2022b).
In its current implementation, FROSch assumes a one-to-one correspondence of subdomains and MPI ranks. Using an interface to the other solver packages in Trilinos, a variety of inexact solvers can be employed for the solution of the local subdomain problems. An extension to multiple subdomains per MPI rank is currently being implemented. Using Kokkos and Kokkos Kernels, FROSch has recently also been ported to GPUs (Yamazaki et al., 2023) with performance gains for the triangular solve or inexact solves with ILU on GPUs; speedups of up to $4.8\times$ are observed when comparing ILU solve times on CPU and GPU.
A demo/tutorial for FROSch can be found at the GitHub repository (45).
@@ 3.4. Multigrid Methods: MueLu
MueLu is a flexible and scalable high-performance multigrid solver library. It provides a variety of multigrid algorithms for problems ranging from Poisson-like operators, over elasticity, convection-diffusion, Navier-Stokes, and Maxwell’s equations, all the way to multigrid methods for coupled multiphysics systems. Besides its strong focus on aggregation-based algebraic multigrid (AMG) methods, MueLu comes with specialized capabilities for (semi-)structured grids to perform semi-coarsening along grid lines while forming the coarse operator via a Galerkin product (in contrast to classical geometric multigrid methods). MueLu is easily extensible which makes it a good framework for the research and development of new multigrid preconditioning methods, e.g. different types of strength-of-connection or drop schemes. Its weak and strong scalability up to $131,000$ cores of a Cray XC40 and one million cores of a Blue Gene/Q system, even for vector-valued partial differential equations (PDEs) on unstructured meshes, have been shown in (Lin et al., 2017; Thomas et al., 2019).
@@
  • Multigrid for multiphysics problems: MueLu implements a tool box to compile multi-level block preconditioners for block matrices arising from coupled multiphysics problems. Applications range from Navier-Stokes equations over surface-coupling (as in fluid-structure interaction or contact mechanics (Wiesner et al., 2021)) to volume-coupled problems (e.g., in magneto-hydro dynamics (Ohm et al., 2022)).
@@
There are a variety of resources to learn more about MueLu. An overview is given on the MueLu website [^12] https://trilinos.github.io/muelu.html. The MueLu User’s Guide (Berger-Vergiat et al., 2023) summarizes installation instructions and a reference to most of MueLu’s configuration parameters. The MueLu Tutorial (Mayr et al., 2025) introduces beginners and experts to various topics in MueLu and shows how to solve or precondition different linear systems using MueLu. Details on the compatibility of MueLu and its predecessor ML (Heroux et al., 2005; Gee et al., 2006) can be found in the MueLu User’s Guide (Berger-Vergiat et al., 2023).
Besides its C++ API, MueLu offers a MATLAB interface, MueMex, to provide access to MueLu’s aggregation and solver routines from MATLAB. MueMex allows users to setup and solve arbitrarily many problems with either MueLu as a preconditioner, Belos as a solver and Tpetra for data structures.
@@
3.5. Direct Linear Solvers: Amesos2 and ShyLU
Amesos2 (Bavier et al., 2012) serves as the central collection of direct solvers for any type of linear systems: it natively implements the direct solver KLU2 (Davis and Palamadai Natarajan, 2010), it offers a uniform interface to third-party direct linear solvers such as CHOLMOD, MUMPS, Pardiso_MKL, SuperLU, SuperLU_MT, SuperLU_Dist, and STRUMPACK, and it provides an interface to the on-node sparse direct solvers implemented in ShyLU.
ShyLU (Thornquist and Rajamanickam, 2014) provides an implementation of three solvers: Basker, Tacho, and a distributed-memory solver based on the Schur complement method, which all have been developed for specific target applications.
Basker (Booth et al., 2017) is a shared-memory sparse direct solver based on LU factorization for the problems that have the block triangular form (BTF) typically seen in circuit simulation applications. Basker uses this structure to factor and solve the diagonal blocks in parallel. The larger diagonal blocks can themselves be factored in parallel by discovering the parallelism available using a nested-dissection reordering. Basker focuses on exploiting thread-parallelism on the multi-core CPU architectures. Amesos2 also has a templated implementation of a sequential KLU solver called KLU2, which also exploits the BTF structure.
Tacho is a shared-memory sparse direct solver that exploits supernodal block structure commonly found in sparse direct factorizations of matrices from mechanics applications. Tacho exploits this supernodal structure for both factorization and triangular-solve phases. It is based on Kokkos. Originally, Tacho implemented a task-parallel Cholesky factorization of sparse symmetric positive definite (SPD) matrices (Kim et al., 2018). However, to improve its portability, it has been extended to compute the sparse factorization based on level-set scheduling. Moreover, its functionality has been extended to the computation of an LDLT factorization of symmetric indefinite matrices, as well as the computation of an LU factorization of general matrices with a symmetric sparsity pattern.
@@ 3.6. Physics Block Operators and Preconditioners: Teko
Many multi-physics problems are conveniently described by a block-structured matrix which allows easy access to the underlying matrices representing single-field physics or coupling between different physical fields. The Teko library (Cyr et al., 2016) provides interfaces for operators and preconditioners that are constructed from such large physics-based sub-blocks. These sub-blocks are Thyra operators which are themselves often implemented using Tpetra matrices.
Teko provides generic block preconditioning strategies such as $2\times 2$ block-LU factorizations as well as block versions of Jacobi- and Gauss–Seidel-type methods for any number of blocks, as well as commonly used approximate inverse strategies for the Navier-Stokes equation such as SIMPLEC, LSC, and PCD preconditioners (Cyr et al., 2012). The Block-LU class offers the possibility to implement user-specific inverse approximation strategies, for example using a sparse approximate inverse to construct the Schur complement (Firmbach et al., 2024). More complicated multilevel hierarchies of block solvers can be generated via Teuchos::ParameterList objects. Block preconditioners for first-order formulations of Maxwell’s equations and Darcy flow are implemented in the Panzer package. Teko’s solvers can be registered in the Stratimikos interface for usage entirely driven by Teuchos::ParameterList objects.
@@ 3.8. Unified Solver Interface: Stratimikos
The Stratimikos package offers a cohesive interface to various linear solvers and preconditioners within Trilinos, including Amesos2, Belos, FROSch, Ifpack2, MueLu, and Teko. It requires that the matrix, right-hand side, and the solution vectors all support the Thyra interface. Thyra also provides wrappers for Tpetra linear algebra. Users can specify solver and preconditioner parameters via a Teuchos::ParameterList, which can be easily populated from an XML file. Moreover, the package allows for the easy addition of new solver and preconditioner factories via adapters, with examples available in the packages/stratimikos/adapters directory. Stratimikos does not support eigenvalue problems. Anasazi has, however, a Thyra interface and can therefore be used with different linear algebra backends.
@@ 4. Nonlinear Solvers and Analysis Tools
The packages included in this section provide the top-level algorithms for computational simulations and design studies. These include nonlinear solvers, bifurcation tracking, stability analysis, parameter continuation, optimization, and uncertainty quantification. A common theme of this collection is the philosophy of “analysis beyond simulation,” which aims to automate many computational tasks that are often performed by application code users by trial-and-error or repeated simulation. Tasks that can be automated include performing parameter studies, sensitivity analysis, calibration, optimization, and the task of locating instabilities. Additional utilities for nonlinear analysis include abstraction layers and interfaces for application callbacks, as well as automatic differentiation tools that can provide the derivatives critical to analysis algorithms.